Find the charge of a point charge q if its electric field E has 12.1 N/C and a distance of 1.8m


1
Expert's answer
2022-05-23T10:18:13-0400
"E=k\frac{q}{r^2}"

"q=Er^2/k\\
=12.1*1.8^2/9*10^9=4.4*10^{-9}\:\rm C"